# Explanation Regularisation through the Lens of Attributions

### v1

@@ 5 Experimental Setting
Table 1: F1-Macro scores ($\uparrow$) and standard deviation for in-domain (ID) and out-of-domain (OOD) data. Attention corresponds to a model that uses attention as the guided attribution technique, and +Rollout to a model that uses attention-rollout. Results are averages of 15 seeds.
Figure 2: F1-Macro scores ($\uparrow$) for both in-domain and out-of-domain data. ER-A is an ER model that uses attention as the guided attribution technique, and ER-R is an ER model that uses attention-rollout. The C prefix indicates a constrained model. Results correspond to 15 seeds.
5.1 Attribution Techniques
We evaluate the plausibility of attributions using four post-hoc attribution techniques to compute token-level global attributions: InputXGradient (IxG; Shrikumar et al., 2017), a gradient-based method that obtains attributions by multiplying inputs by their respective gradients; ALTI Ferrando et al. (2022), a vector-based technique that measures token-to-token interactions at a layer level,and that linearly combines these interactions to provide global attribution scores; and DecompX (DX; Modarressi et al., 2023), a global vector-based technique based on the propagation of decomposed token representations through the model (we also use a variant, DX-C, that includes the classification head). For IxG we use the Captum package (Kokhlikyan et al., 2020). For DX, DX-C [^4] https://github.com/mohsenfayyaz/DecompX and ALTI [^5] https://github.com/mt-upc/transformer-contributions we adapt the available code.
5.2 Plausibility Metrics
To measure how well the post-hoc input attributions reproduce the human annotated rationales, we employ three metrics introduced by Fomicheva et al. (2021), and further described in Appendix B: AUC Score, Average Precision; and Recall@k.
6 Results
6.1 Classification Task Generalisation
We start by reporting how ER+Att classification performance generalises to OOD data. Table 1 shows the average F1-Macro scores for both in- and OOD data. Similarly to Joshi et al. (2022) and Stacey et al. (2022), we observe a positive impact in OOD average classification performance when training with ER. However, if we consider standard deviation, and visualise the distribution of results for all model seeds (Figure 2), it is noticeable how the effect on the reported average improvements are exacerbated by certain model versions that perform poorly. In fact, this seems to indicate that one of the merits of ER is to converge less often to models that generalise poorly to OOD conditions.
6.2 ER Impact on Model Attributions
Existing work on explanation regularisation has shown evidence of improved alignment between the output of the guided attribution technique and human annotated rationales (Stacey et al., 2022; Madani and Minervini, 2023). However, this analysis is limited to the guided attribution technique (i.e., the one used during training). We argue that such improvements are to be expected, provided that the model converges to a solution that minimizes the explanation loss. However, attributions revealed by an attribution technique are not necessarily aligned with the model’s decision process—there exists a faithfulness gap for input attribution methods. The increased alignment between the guided attention technique and the human rationale may have been achieved by broadening the gap rather than making the ER model more reliant on plausible tokens. In order to test this hypothesis, we compute plausibility metrics (Section 5.2) for attributions with a series of techniques described in Section 5.1. If ER genuinely improves model reliance on plausible features then we would expect to observe changes in attributions for other techniques, rather than only for the one used in training. We show average scores for AUC in Table 2.
As expected, ER+Att is able to strongly impact the average AUC plausibility of the guided attribution technique (Att), by increasing its value from 35.6 (in the baseline) to 78.6. However, for the remaining suite of attribution techniques, the impact is small or non-existent (e.g., 50.6 vs. 51.5 for IxG). Thus, we can indeed conclude that the improvement in plausibility for attribution obtained with Att was achieved largely by broadening the faithfulness gap rather than by making the ER model more reliant on plausible features. [^6] We find similar evidence using average precision and recall@k plausibility metrics (Appendix Table 8).
@@ 6.2 ER Impact on Model Attributions
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$
@@
Table 2: Average AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A) and $\mathcal{L}_{\text{expl}}$ (R) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att and AttR, respectively. Highlighted values correspond to the guided attribution technique used for ER.
6.3 ER with Global Attribution Technique
We could hypothesise that the failure to enhance the genuine reliance of the method on plausible features is due to the use of attention in ER. It is evident that top-layer attention can align with human rationales without altering the underlying function, for example, by rescaling its input vectors to compensate for the change in attention or by shifting ‘implausible’ computations to a lower layer, thus, hiding them from the regulariser. Therefore, here, we test whether the same behavior is observed with a more reliable global attribution technique. We choose attention-rollout (AttR; Abnar and Zuidema, 2020), a computationally efficient way of obtaining differentiable global vector-based attributions (see Section 4.1).
In terms of classification performance, ER+ AttR performs similarly to ER+Att (Table 1). Also, despite its potential to lead ER models to rely on plausible features, we find once again that other attributions are not affected (Table 2). This seems to indicate that a joint loss setup in ER does not result in models that lead to more plausible attributions overall. It is also noticeable how, contrarily to what we have seen with ER+Att in Section 6.2, ER+AttR almost did not impact the respective guided attribution technique (28.7 vs 27.9). To understand why this might have happened, we train ER models exclusively with the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, and obtain ‘optimal’ loss values. We then compare these with the values obtained for the joint ER models. As observed in Table 3, there is a gap to the ‘optimal’ explanation loss value for both approaches, more so for AttR, extending to AUC—for $\mathcal{L}_{\text{expl}}$ (A) we obtain a AUC value of 89 (vs the 78.6 obtained with joint ER+Att), and for $\mathcal{L}_{\text{expl}}$ (R) we obtain a value of 89.4 (vs the 28.7 obtained with joint ER+AttR). This shows that ER models optimised with a joint objective can converge to a solution where the model effectively ignores the auxiliary explanation loss component.
Att $\mathcal{L}_{\text{expl}}$ 0.030 89.0
(J) ER 0.040 78.6
(C) ER 0.034 87.6
AttR $\mathcal{L}_{\text{expl}}$ 0.032 89.4
(J) ER 0.064 28.7
(C) ER 0.033 89.1
Table 3: SST-Dev explanation loss and AUC values for models guided with attention (Att) or attention-rollout (AttR), as a function of their training objective (explanation loss alone, vs. (J)oint ER and (C)onstrained ER).
6.4 Constrained Optimization Approach
We have observed how the joint objective might result in ER models that ‘under-optimise’ the explanation loss. Yet, the results shown in Table 2 for the models trained exclusively with the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, show an interesting observation: attention-rollout as the guided attribution technique leads to a model that more strongly relies on plausible tokens, as observed by the positive impact across all reported attribution techniques. In order to bridge the gap between these two observations, we make use of constrained optimisation (Section 4.2). The goal is that, by constraining ER to converge to solutions that are close to the ‘optimal’ explanation loss, while still optimising for classification performance, we can better evaluate the impact of using models whose classification decisions rely more on plausible tokens.
As observed in Table 2, while constrained ER+Att still mostly fails to impact the evaluation attribution techniques, constrained ER+AttR impacts the plausibility of all assessed techniques. However, despite the positive impact of constrained ER+AttR on the plausibility of attributions, Table 1 indicates a decline in classification performance. In summary, the only ER strategy we tested that results in a model relying more strongly on plausible tokens decreases OOD classification robustness. This suggests a disconnect between (over-)reliance on plausible features and OOD robustness.
6.5 Classification/Explanation Loss Trade-off
Figure 3: Cross-Entropy vs Explanation Loss for SST-Dev. The vertical lines show the explanation-loss bound. Each point corresponds to the average of 5 runs, and the error bars to the standard deviation for each loss. $\lambda$ ranges from 0 (No-ER) to 100.
We have observed how constrained optimisation leads to models that are able to retain some of the classification task performance, but do not improve the plausibility of input attributions (Constrained ER+Att), or models that do improve the plausibility of attributions, but at the expense of the classification task performance (Constrained ER+AttR). In order to better understand how the choice of ER attribution technique impacts the trade-off between the two losses we are jointly optimising, we study this scenario with increasing values of $\lambda$.
We observe two overall trends for SST-Dev data (Figure 3). Given a high enough $\lambda$, the explanation loss converges to the ‘optimal’ empirical value. For ER+Att, smaller values of $\lambda$ are enough to impact the explanation loss, whereas for ER+AttR larger values are required. For both strategies alike, as the explanation loss values decrease, the cross entropy loss degrades (we do verify that all models are still learning to solve the classification task). However, the ER+Att model seems to be more robust to this trade-off—we observe a smaller drop in CE loss as the explanation loss decreases when compared to ER+AttR. This is aligned with our previous observation: improving the explanation loss (i.e. the model is better at predicting the human rationales) negatively impacts the CE loss, decreasing classification task performance, where ER+Att outperforms ER+AttR (Table 1).
The presence and the prominence of this trade-off may seem puzzling. We hypothesise two potential, non-mutually exclusive explanations. First, human rationales may not accurately reflect features useful for models due to noise in annotation or the inability of annotators to identify all genuinely predictive features (Carton et al., 2020). Second, attention regularisation imposes assumptions about how the model should implement the function (e.g., aggregating information in the top layer for ER+Att), which may be sub-optimal for the task. Investigating these hypotheses requires a range of experiments across various tasks and models, which is beyond the scope of this paper. However, we can already conclude that the presence of the trade-off reduces the opportunity to simultaneously improve both OOD performance and reliance on plausible features.
6.6 Predicting OOD Performance
As shown in Figure 2, ER models exhibit a broad spread of OOD classification results, highlighting the importance of identifying the best-performing models, ideally without access to OOD data. With that goal, we investigate whether it is possible to predict model performance, by studying if (i) in-domain classification or plausibility performance predicts OOD classification performance; (ii) OOD plausibility predicts OOD classification performance; and (iii) the auxiliary task transfers OOD.
(a) Predictor: SST-Dev Classification
(b) Predictor: SST-Dev Plausibility AUC w/ DX-C
(c) Predictor: Movies / Yelp-50 Plausibility AUC w/ IxG
From In-Domain Performance.
Figures 4(a) and 4(b) show the association between in-domain (SST-Dev) predictors classification F1-Macro and plausibility AUC (computed with DX-C), and classification F1-Macro for SST-Test (in-domain) and Movies (OOD). In both cases there is no clear correlation between the in-domain metrics and OOD classification scores, with the result in Figure 4(b) further supporting the apparent disconnect between reliance on plausible features and OOD classification performance. We find a similar behavior for other approaches, OOD datasets, plausibility metrics and attribution techniques (Appendix D).
From OOD Plausibility Performance.
Given that we do not find in-domain classification performance or attribution plausibility to be predictive of OOD performance, we now investigate if there is any correlation at the OOD level. In order to test this, we use the Movies dataset, and Yelp-50, a subset of Yelp, annotated as described in Appendix E. Figure 4(c) shows the association between OOD plausibility AUC (computed with IxG [^7] Movies consists of long inputs, resulting in prohibitive runtime for batch sizes that do not result in OOM errors, on a Nvidia A100, when using DX and ALTI. Thus, we report IxG, as it can be computed for both datasets in the figure.) and classification F1-Macro, for both datasets. Similarly to what we observed in-domain, there is no correlation between OOD plausibility scores and its corresponding classification task performance.
ER Impact on OOD Attributions.
If improved OOD classification performance is to be ascribed to more plausible attributions, we would expect ER to impact plausibility metrics OOD. Results for AUC (Table 4) show that the impact of ER on OOD attributions is still present, but more limited than in-domain (Table 2), hinting at a limited OOD impact of the auxiliary task introduced in ER.
@@ 6.6 Predicting OOD Performance
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$
@@
Table 4: Average AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A) and $\mathcal{L}_{\text{expl}}$ (R) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att and AttR, respectively. Highlighted values correspond to the guided attribution technique used for ER.
@@ 7 Conclusion
We find that ER, unless constrained to meet an ‘optimum’ explanation loss computed with a global attribution technique, does not lead to models that effectively attribute to more plausible tokens. In fact, we observe a disconnect between reliance on plausible tokens and OOD robustness, with (i) OOD classification improvements not being predicted by increased plausibility; and (ii) OOD classification performance degrading when models are constrained to rely more on plausible tokens.
Our findings highlight relevant challenges for ER and suggest future research questions: (i) how to achieve a better balance between constraining the model to attend to plausible tokens and solving the classification task, e.g., by learning which layers and attention heads of the Transformer model to regularise (see (Fernandes et al., 2022)), or (ii) by allowing the model to learn to select which rationales to use during training (see (Arous et al., 2021; Carton et al., 2022)); and (iii) how would input attributions for an ER model be impacted by using more faithful, while still efficient, vector-based global guidance techniques as part of ER design choices (and not only as an analysis tool).
@@
Data.
In terms of data, we identify two independent limitations. The first limitation related with data concerns the limited number of datasets per task that include annotations, making it difficult to analyse the impact of ER on the attributions on out-of-domain settings. Secondly, we use only the task of sentiment analysis to study the disconnect between ER’s robustness to OOD conditions and the plausibility of attributions. In practice, it would be possible to conduct a similar analysis for other datasets that include human annotations. However, the questions we ask are deeply inherent to ER, and it seems unlikely that alternative tasks would be impacted differently.

### v3

@@ 5.1 Joint Optimisation
IxG 90.64 $\pm$ 0.90 91.59 $\pm$ 1.46 75.80 $\pm$ 15.1 88.99 $\pm$ 6.17 83.35 $\pm$ 7.58 79.95 $\pm$ 2.57
Table 2: F1-Macro ($\uparrow$) and standard deviation for in-domain (SST) and OOD data. Attention corresponds to an ER (J)oint or (C)onstrained model that uses attention as the guided attribution technique, +Rollout to a model that uses attention-rollout, and IxG to a model that uses InputXGradient. Results are averages of 15 seeds.
Figure 2: F1-Macro scores ($\uparrow$). ER + Att uses attention as the guided attribution technique, ER + AttR uses attention-rollout, and ER + IxG uses InputXGradient. The C prefix indicates a constrained model. Results correspond to 15 seeds. ER-C + IxG is not shown for improved clarity and can be seen in Appendix Figure 11.
Only local guidance improves rationale extraction performance.
In addition to classification, ER models are trained to align guided attributions to human rationales. Hence, compared to the baseline, we expect an impact on the plausibility of the corresponding guided attributions. This effect can be observed by comparing the highlighted cells in Table 3 with the corresponding baseline. Our expectation is met only with local guidance, with average AUC plausibility increasing from 35.6 (baseline) to 78.6 (w/ Att) in-domain and similar, but more modest, improvements OOD (Table 4).
@@
w/ IxG 63.9 42.9 81.5 74.8 61.3 64.8
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$5.1)
@@
Table 3: Average in-domain AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A/R/I) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att, AttR, and IxG respectively. Highlighted values correspond to the ER guided attribution technique.
@@
w/ IxG 56.7 56.1 61.6 60.8 50.5 69.8
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$5.1)
@@
Table 4: Average out-of-domain AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A/R/I) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att, AttR, and IxG respectively. Highlighted values correspond to the ER guided attribution technique.
The lack of impact of global guidance on the plausibility of the corresponding guided attributions might come as a surprise. However, we need to consider that: (i) this class of techniques incorporates more of the model’s components when computing attributions, potentially making it more difficult for the model to ‘hide’ implausible computations from the explanation loss; and (ii) joint ER balances both classification and explanation losses with a preference (via model selection) for classification, meaning that to maintain classification performance the model might require ‘under-optimising’ the explanation loss. Both possibilities will be further studied in this section.
ER+Att fails to impact non-guided attribution techniques.
The previous results illustrate how attention guidance (ER+Att) is able to simultaneously improve OOD classification and rationale extraction (when assessed with the corresponding guided attribution technique). It is tempting to connect these two observations and conclude that Att guidance results in ER models that better classify while relying more on plausible tokens. However, we find that not to be the case. Firstly, we correlate input attributions (obtained by a given technique) across training conditions (e.g., ALTI baseline attributions vs. ALTI ER+Att attributions). We find that – aside from the guided attribution technique – correlation coefficients are mostly unaffected by ER training, both in- and out-of-domain (see first row of Figure 3, for OOD). We find similar evidence for all non-guided techniques (Appendix Fig. 9). Secondly, we compare the AUC plausibility scores of the non-guided techniques—any non-highlighted cell in Tables 3 and 4—with the corresponding baselines. Here, we observe a lack of impact across joint ER models. For example, for ER+Att we find the impact to be small (59.0 vs 64.5 for ALTI) to none (58.5 to 58.4 for DX-C). We find similar evidence using average precision and recall@k (Appendix Table 9). In fact, if we look at AUC plausibility scores across layers (Fig. 4) we can observe that ER+Att only impacts the plausibility of attributions at the top-layer. This result confirms our suspicions: local guidance is able to ‘hide’ implausible computations at lower layers, thus hiding them from $\mathcal{L}_{\mathrm{expl}}$.
Figure 3: Yelp-50 (OOD) Kendall Rank correlations between attribution techniques for different approaches. Baseline (BS) vs Baseline serves as ground-truth for the expected correlations agreement due to seed variability.
Figure 4: SST-Dev average AUC plausibility score per layer ($\uparrow$) with Attention and DecompX.
$\mathcal{L}_{\mathrm{cls}}$ and $\mathcal{L}_{\mathrm{expl}}$ exhibit no synergy.
We observed before how ER+AttR and ER+IxG failed to impact input attributions, including for the respective guided attribution technique. To understand why this happened, we train ER models exclusively with the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, and obtain a ‘lower-bound’ for its value. As seen in Table 5, all joint ER approaches exhibit a gap to the explanation loss ‘lower-bound’ value, particularly AttR and IxG, with this observation extending to AUC — e.g., for $\mathcal{L}_{\text{expl}}$ (A) we obtain a AUC value of 89.0 (vs the 78.6 obtained with joint ER+Att), and for $\mathcal{L}_{\text{expl}}$ (R) we obtain a value of 89.4 (vs the 28.7 obtained with joint ER+AttR).
Att $\mathcal{L}_{\text{expl}}$ 0.030 89.0
(J) ER 0.040 78.6
(C) ER 0.034 87.6
AttR $\mathcal{L}_{\text{expl}}$ 0.032 89.4
(J) ER 0.064 28.7
(C) ER 0.033 89.1
IxG $\mathcal{L}_{\text{expl}}$ 0.349 84.3
(J) ER 0.432 53.8
(C) ER 0.358 81.5
Table 5: SST-Dev explanation loss and AUC values for models guided with Att, AttR, or IxG, as a function of their training objective (explanation loss alone, vs. (J)oint ER and (C)onstrained ER).
These results highlight one possible difficulty of the joint ER setup: promoting plausible attributions via ER with global guidance might incur a negative cost in classification performance, hence joint optimisation (whose hyperparameters are selected based on classification loss) will under-optimise the auxiliary task. In order to confirm this behavior we run the joint ER setup with increasing values of $\lambda$. The results for ER+Att and ER+AttR can be seen in Fig. 5. We observe two overall trends. Given a high enough $\lambda$, the explanation loss converges to the obtained ‘lower-bound’. For ER+Att, even smaller values of $\lambda$ impact the explanation loss, whereas for ER+AttR larger values are required. For both strategies alike, as the explanation loss values decreases, the cross entropy loss increases (we verify that all models still converge on the classification loss). However, the ER+Att model seems to be more robust to this trade-off—we observe a smaller increase in CE loss as the explanation loss decreases when compared to ER+AttR. We observe a similar effect for ER+IxG in Appendix Fig. 12.
Although this is, perhaps, not a necessarily surprising finding (Carton et al. (2022), for example, observe a similar tradeoff, though in a multi-tasking setting), it challenges a key assumption of ER as a means to improve OOD generalisation: that the explanation loss should help inform the classifier, improving OOD performance. In fact, we observe that the joint ER setup will struggle to accommodate both losses, particularly when using a guided attribution technique that is more difficult to ‘hack’ in order to become apparently better at predicting human rationales, such as attention-rollout, where the explanation loss might be lowered, but far from optimally.
Figure 5: SST-Dev $\mathcal{L}_{\text{ce}}$ vs. $\mathcal{L}_{\text{expl}}$. The vertical lines show the validation explanation-loss bounds. Each point is the average of 5 runs, the error bars show standard deviation for each loss. $\lambda$ ranges from 0 (No-ER) to 100.
5.2 Constrained Optimisation
In §5.1, we observed how joint ER under-optimises the explanation loss and how this effect varies in function of the guided attribution technique being local or global. Moreover, Table 3 shows that optimising exclusively the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, can lead to AUC plausibility improvements—on the guided attribution technique alone for local guidance, and across the board for global guidance. This serves as motivation to analyse classifiers optimised for the classification task, while subject to achieving the aforementioned explanation loss ‘lower-bound’ via constrained optimisation.
Constrained global guidance is necessary to impact non-guided attribution techniques.
We start by noting how constrained ER addresses an earlier result: both global guided techniques are now impacting the corresponding guided attribution technique. This is true for both the AUC plausibility scores (in Table 3) and correlation coefficients (in Appendix Figure 9). However, this is so by design, and a more interesting observation stems from inspecting the non-guided attribution techniques — while constrained ER-C+Att still mostly fails to impact other attribution techniques, constrained ER-C+AttR and ER-C+IxG impact the outcome of all assessed attribution techniques. We can observe this along three ‘dimensions’. First, global techniques are able to clearly impact the plausibility of all attribution techniques (Table 3). For instance, ER-C+AttR leads to large improvements in DecompX, where the average AUC plausibility score is improved from 57.5 (in baseline) to 81.3, with the value being 56.8 for ER-C+Att. Second, there is a clear impact in the correlation of attributions across approaches (Appendix Fig. 9(a)), contrarily to what is observed for local ER, including ER-C+Att. Finally, AUC plausibility scores per layer (Figure 4) also show how the global technique is able to impact not only attributions at the top-layer, but across the whole model, making them more aligned with the human rationales. All observations along the three inspected ‘dimensions’ hold for the OOD data, with attributions clearly being impacted (Figure 3), despite the lower influence in the AUC plausibility metric (Table 4). [^3] We also find similar evidence for AUC per-layer across multiple attribution techniques (Appendix Figure 13).
There is a disconnect between (over-)reliance on plausible features and OOD robustness.
Despite the noticeable impact of constrained ER-C+AttR and ER-C+IxG on attributions and their plausibility, Table 2 indicates a decline in classification performance. That is, the only ER strategies that resulted in models that rely more strongly on plausible tokens decrease OOD classification robustness. This is somewhat expected given what we observed before in Figure 5. However, it does lead us to conclude that the ER protocol should be re-considered, or at least, a bigger emphasis on trying to understand what ‘amount’ of increased plausibility is desirable, and where that information should be used while regularising the model, in order to better align the current expectations of improved OOD generalisation due to an increased reliance of ER models on plausible features.
5.3 Predicting OOD performance
As shown in Figure 2, ER models exhibit a broad spread of OOD classification results. This highlights the importance of identifying the best-performing models, ideally without access to OOD data. Thus, we investigate whether it is possible to predict OOD classification performance from features potentially available for model selection.
(a) ID Classification
(b) ID Plaus. AUC w/ DX-C
(c) OOD Plausibility AUC w/ IxG
In-domain classification and plausibility performance are not predictive of OOD improvements.
Figures 6(a) and 6(b) show the association between in-domain (SST-Dev) predictors classification F1-Macro and plausibility AUC (computed with DX-C), and classification F1-Macro for Movies (OOD). In both cases there is no clear correlation between the in-domain metrics and OOD classification scores, with the result in Figure 6(b) further supporting the apparent disconnect between reliance on plausible features and OOD classification performance. We find a similar behavior for other approaches, OOD datasets, plausibility metrics and attribution techniques (Appendix F).
OOD plausibility performance is not predictive of OOD improvements.
Given that we do not find in-domain classification performance or attribution plausibility to be predictive of OOD performance, we now investigate if there is any correlation at the OOD level. In order to test this, we use the Movies and Yelp-50 datasets. Figure 6(c) shows the association between OOD plausibility AUC (computed with IxG [^4] Movies consists of long inputs, resulting in prohibitive runtime for batch sizes that do not result in OOM errors, on a Nvidia A100, when using DX and ALTI. Thus, we report IxG, as it can be computed for both datasets in the figure.) and classification F1-Macro, for both datasets. Similarly to what we observed in-domain, there is no correlation between OOD plausibility scores and its corresponding classification task performance.
The combination of both these results highlights the difficulty of predicting OOD results, and brings forth another challenge of the ER setup, where we need to make decisions about design choices based on in-domain performance, that does not seem to be linked to OOD generalisation.
@@ 6 Conclusion
We find that ER, unless constrained to meet a ‘lower-bound’ on the explanation loss computed with a global attribution technique, does not lead to models that effectively rely on plausible tokens. In fact, we observe a disconnect between reliance on plausible tokens and OOD robustness, with (i) OOD classification improvements not being predicted by increased plausibility; and (ii) OOD classification performance degrading when models are constrained to rely more on plausible tokens.
Our findings highlight relevant challenges for ER and suggest future research questions: (i) how to achieve a better balance between solving the classification task and constraining the model to attend to plausible tokens, e.g., by learning which layers and attention heads of the Transformer model to regularise (see Fernandes et al. (2022)), or (ii) by allowing the model to learn to select which rationales to use during training (see Arous et al. (2021); Carton et al. (2022)); and (iii) how would input attributions for an ER model be impacted by using more informed, while still efficient, vector-based global guidance techniques as part of ER design choices (and not only as an analysis tool).
@@
Data.
In terms of data, we identify two limitations. The first limitation related with data concerns the limited number of datasets per task that include annotations, making it difficult to analyse the impact of ER on the attributions on out-of-domain settings. Secondly, we use only the task of sentiment analysis to study the disconnect between ER’s robustness to OOD conditions and the plausibility of attributions. This is also connected to the first discussed limitation — a full ER setup requires rationale-annotated data of an appreciable dimension for training, and also multiple OOD datasets, including at least one with rationale-annotated data. However, the questions we ask are deeply inherent to ER, and it seems unlikely that alternative tasks would be impacted differently.